What is the energy of microwave radiation that has a wavelength of 0.292 m?
Answer #1

А= 0. 2-12 м h= 6-62 гло Чтке с = 3x to m| к Е. 66 25 по 24 тяx зxo kxt о: 27. L) - 68. 06 х 10 Т Г 10 -. С. Ro6X (о т Е. С.